Which of the following is equal to 4/7*14/3? A. 4/7*3/14 b.4*2/3 c.7/4*14/3 d. 2/7*7/3
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Which of the following is equal to
4/7*14/3 = (4/1)(2/3) = 8/3 = 2 2/3
-----------------------------------
A. 4/7*3/14 = 12/98 :: NOT
--------
b.4*2/3 = 8/3 = 2 2/3 :: Yes
--------------------------------
c.7/4*14/3 = 98/12:: NOT
--------------------------------
d. 2/7*7/3 = 2/3:: NOT
============================
cheers,
Stan H.
=======================

Answer by ewatrrr(24785)